# Adaptive Pruning of Pretrained Transformer via Differential Inclusions

### v1

@@ 5.1 Main results
Table 3: Ablation: The results of CLIP-large and CLIP-base.
Transformer weight family results of DeiT. Table 1 shows our transformer weight family results of DeiT on ImageNet-1k. The results show that our proposed method performed well on DeiT. For DeiT-Small, our method maintained accuracy at 80.2% while reducing 29.5% of the parameters. Compared to the other methods, WDpruning only achieved 78.6% accuracy with a 29.6% compression ratio. Only UPop Shi et al. (2023) achieved comparable results to ours. This indicates that our method of solution path is highly effective, preserving more parameter information even at higher compression ratio.
Our method performed well on DeiT-Base. At almost the same compression ratio, our accuracy surpassed SCOP Tang et al. (2020) and PoWER Goyal et al. (2020), and we achieved a better compression ratio than IA-RED  Pan et al. (2021), the smallest model among the compared algorithms.
@@
Results of tiny models and datasets. We also extended our method to the Tiny models like Swin-Tiny and DeiT-Tiny model. Table 2 presents a comparison between our approach and the other method, ViT-Slim Chavan et al. (2022). Notably, our method achieves only a 0.8% reduction in accuracy while compressing 5% more than ViT-Slim Chavan et al. (2022). The success of our method with the Swin-T model, which demonstrates its adaptability across various transformer models.
For DeiT-Tiny, at almost the same compression ratio, our accuracy surpassed SSP Chen et al. (2021) and $\text{S}^{\mathrm{2}}$iTE Chen et al. (2021) by around 2% points, and we achieved a better compression ratio than GOHSP Yin et al. (2023), the smallest model among the compared algorithms.
@@ 5.3 Further studies
We demonstrated the consistency of the transformer weight family through experiments and then applied SPP to LLMs.
@@
Extentions to LLMs. We extended our method to large language models (LLMs) with post-training. We applied our solution path method during the LLM pruning search stage, combining it with the RIA  Zhang et al. (2024) pruning metric. The detailed algorithm is shown in Algorithm 2 .
We applied our method on Llama2-7B and OPT-6.7b. The calibration datasets C4 and WikiText2 were used to generate activations during the forward pass, which, along with weight magnitude, served as the pruning metric. The results were reported on 5 datasets. As shown in Table 5, the accuracy after pruning with 50% sparsity is comparable to the advanced method RIA and Wanda Sun et al. (2023), indicating the potential our method in pruning LLMs.

### v2

@@ 5.1 Main results
Table 3: Ablation: The results of CLIP-large and CLIP-base.
@@
Transformer weight family results of DeiT. Table 1 shows our transformer weight family results of DeiT on ImageNet-1k. The results show that our proposed method performed well on DeiT. For DeiT-Small, our method maintained accuracy at 80.2% while reducing 29.5% of the parameters. Compared to the other methods, WDpruning only achieved 78.6% accuracy with a 29.6% compression ratio. Only UPop (Shi et al., 2023) achieved comparable results to ours. This indicates that our method of solution path is highly effective, preserving more parameter information even at higher compression ratio.
Our method performed well on DeiT-Base. At almost the same compression ratio, our accuracy surpassed SCOP (Tang et al., 2020) and PoWER (Goyal et al., 2020), and we achieved a better compression ratio than IA-RED  (Pan et al., 2021), the smallest model among the compared algorithms.
@@
Results of tiny models and datasets. We also extended our method to the Tiny models like Swin-Tiny and DeiT-Tiny model. Table 2 presents a comparison between our approach and the other method, ViT-Slim (Chavan et al., 2022). Notably, our method achieves only a 0.8% reduction in accuracy while compressing 5% more than ViT-Slim (Chavan et al., 2022). The success of our method with the Swin-T model, which demonstrates its adaptability across various transformer models.
For DeiT-Tiny, at almost the same compression ratio, our accuracy surpassed SSP (Chen et al., 2021) and $\text{S}^{\mathrm{2}}$iTE (Chen et al., 2021) by around 2% points, and we achieved a better compression ratio than GOHSP (Yin et al., 2023), the smallest model among the compared algorithms.
@@ 5.3 Further studies
Extentions to LLMs. We extended our method to large language models (LLMs) with post-training. We applied our solution path method during the LLM pruning search stage, combining it with the RIA  (Zhang et al., 2024) pruning metric. The detailed algorithm is shown in Algorithm 2 .
We applied our method on Llama2-7B and OPT-6.7b. The calibration datasets C4 and Wikitext2 were used to generate activations during the forward pass, which, along with weight magnitude, served as the pruning metric. The results were reported on 5 datasets. As shown in Table 5, the accuracy after pruning with 50% sparsity is comparable to the advanced method RIA and Wanda (Sun et al., 2023), indicating the potential our method in pruning LLMs.
@@ 6 Conclusion
ACKNOWLEDGEMENTS
This work was supported by the Science and Technology Commission of Shanghai Municipality(No. 24511103100).